# Scalable Link Prediction on Large-Scale Heterogeneous Graphs with Large Language Models

### 24 Jan 2024 v1

@@ 2.5 Self-supervised Learning based Finetune
As a more relevant graph structure, large-scale heterogeneous graphs lack labelled data. LPNL supports self-supervised learning based finetune for large language models without relying on labels. During the end-to-end prompt tuning, it automatically constructs a candidate set containing ground truth, aligned with downstream prediction formats. And the ground truth is used as the correct answer for link prediction. To ensure training correctness, the ground truth appears randomly within the candidate sequence, and notably, during the heterogeneous subgraph sampling process in Section 2.3, the edges between the ground truth and the source node are masked.
@@ 3.1 Experiment Settings
Models We finetune Flan-T5 Chung et al. (2022) as the backbone language model for our LPNL. The numbers of sampling hops $h=3$, top anchor nodes sequence $k=50$ and candidate length limit $\mathcal{L}=3$, mentioned in Section 2.3 and 2.4 is used for all LPNL.
Datasets We conducted all experiments on the OAG dataset Huang et al. (2020), known as one of the largest publicly available heterogeneous graphs, comprising 178 million nodes and 2.236 billion edges. It includes five types of nodes (denoted as papers (P), authors (A), venues (V), institutes (I) and fields (F)) and their interrelations. In our specific experiments, we utilized four representative domain-specific subgraphs from OAG: Computer Science (CS), Material Science (Mater), Engineering (Engin) and Chemistry (Chem) Jiang et al. (2021). The graph statistics are listed in Table 1. We partition each dataset into finetune, validation, and test sets based on distinct time periods. Specifically, in the OAG dataset, papers are published between 1900 and 2019. Consequently, we utilize publications preceding 2015 for finetune, data from 2015 to 2016 for validation, and information from 2016 onwards for testing.
Task We consider the real-world link prediction tasks to evaluate the performence of our LPNL: author name disambiguation Ferreira et al. (2012), which is a fundamental challenge for curating academic publication and author information, as duplicated names widely exist in our lives. Specifically, given a paper, the objective is to predict the true author who has a genuine link with that paper among all authors with the same name.
@@ 3.2 Overall Performance
In this experiment, we compare the T5 model as the backbone version of LPNL to advanced GNN based baseline models across the four domain-specific subgraphs. We finetune the model separately across various subgraphs and tasks and evaluate the performance of the models in link prediction. The experimental results of the proposed method and baselines are summarized in Table 2. All experiments for the author name disambiguation task over all datasets are evaluated in terms of NDCG, MRR and Hits@1 Li (2022); Liu et al. (2009).
@@
Figure 5: Cross-Domain Transfer Results
@@ 3.3 Cross-Domain Knowledge Transfer
To explore the generalization capabilities of LPNL, we set up experiments for cross-domain knowledge transfer. Specifically, we fine-tuned the T5 model using LPNL on a subgraph corresponding to one domain and subsequently conducted testing on subgraphs from other domains. The experimental outcomes, visualized in Figure 5 as a heatmap, reveal that in most instances, the model exhibits optimal performance when fine-tuned within its original domain. Surprisingly, models fine-tuned on other domains also demonstrate remarkably strong performance, often closely matching or even surpassing the best performance achieved by fine-tuning within the original domain (e.g., Mater-Engin). This highlights the robust knowledge transferability of our approach, showcasing minimal sensitivity to domain-specific information and demonstrating robust generalization capabilities.
Figure 6: Comparison of few-shot results
3.4 Few shot cability
The extensive pretraining of large language models across various natural language tasks has endowed them with robust reasoning and generalization capabilities. In contrast to traditional GNN models, they require minimal training samples to converge and exhibit superior performance. We further investigates the few-shot capabilities of LPNL by comparing it with the top-performing homogeneous GNN and heterogeneous GNN in terms of overall performance. Configure the evaluate results to be printed every 1 batch, with each batch consisting of 50 link prediction tasks. Compare the few-shot results for the first 20 batches. The results in Table 6 demonstrate that our LPNL swiftly converges with minimal sample fine-tuning, displaying comparable performance to the best fine-tuning outcomes. This showcases the portability of large language models in addressing graph-related tasks.
@@ 3.4 Few shot cability
Table 3: Ablation Study Results of Sampling Methods
@@ 3.5 Ablation Study
We conducte an ablation study on CS dataset to evaluate the effectiveness of our approach in employing large language models combined with graph knowledge strategies. We compare the performance among different versions of sampling methods: the standard LPNL, a version without any graph information, and another two sampling versions, each independently utilizing distinct stages. As illustrated in Table 3, the model’s performance significantly diminishes when graph information is excluded. Furthermore, the performance of the versions without stage 1 and stage 2 shows a notable gap compared to LPNL. It indicates that employing our designed two-stage sampling pipeline enables LPNL to capture crucial information within the graph after balanced heterogeneous sampling, resulting in improved predictive outcomes.
 Hop Info NDCG MRR Hits@1
Multi-hop 97.86 94.37 89.47
@@
Table 4: Ablation Study Results of Multi-hop
In our experiments, two critical operations contributing significantly to the outstanding performance of LPNL in link prediction are multi-hop and anchor nodes, which provide essential information to the LLMs. To assess the impact of these two key components on model performance, we conducted another ablation experiments, and the results are presented in Table 4 and 5. It shows that incorporating multi-hop and more anchor nodes information can both enhance the LPNL’s performance. However, further experiments indicate that increasing the number of multi-hop and anchor nodes beyond a certain threshold does not lead to significant performance improvement. On the contrary, it may result in additional costs without notable benefits.
 Top-k Info NDCG MRR Hits@1
@@
Table 5: Ablation Study Results of Anchor Nodes
@@ 4 Related Work
Large Language Models with Graph Knowledge The emergence of large language models (PLMs) has propelled natural language processing (NLP) to new heights. Qiu et al. (2020). For example, models like BERT Devlin et al. (2018) and T5 Raffel et al. (2020) demonstrate excellent performance in a wide range of downstream tasks, such as text classification and question answering. Besides, some works  Zhang et al. (2019); Liu et al. (2022, 2020)attempts to inject external graph knowledge into LLMs, thus enabling LLMs to gain the ability to solve problems on graphs Recently, due to the powerful inferential capabilities of large language models, a burgeoning body of work Fatemi et al. (2023); Ye et al. (2023); Liu et al. (2023) attempt to utilize natural language descriptions of graph features, employing generated prompts to instruct large language models in addressing various problems on graphs.
@@ 5 Discussion
From our experiments, we found that describing graphs using natural language does not follow the principle of ’more information is better.’ Sampling more nodes can introduce additional information, but it may lead to information redundancy, resulting in a decline in the inferential capabilities of large models. Therefore, the key lies in the setting of the sampling and divide-and-conquer length limits, which should align with the input window size and inferential capabilities of the large language model. While designing prompts, we observe that complex relationships between nodes are challenging to articulate in text, especially in large or dense graphs, potentially leading to redundant contexts. LPNL leverages structural information during the sampling phase and, in the prompt generation, only conveys information about the sampled nodes. This approach aims to minimize context redundancy while maximizing the utilization of graph information.
@@ 6 Conclusion
In this paper, we explore, for the first time, the application of large language models to address the link prediction task on large-scale heterogeneous graphs. We introduce LPNL, a large language model based framework for scalable link prediction on large-scale graphs. We design specific prompt templates for the link prediction task and generate the prompts based on anchor nodes obtained through a two-stage sampling approach. These prompts are then input to the large language model for predictions. To tackle the token overload issue arising from an excessive number of candidate nodes, we employ a divide-and-conquer strategy. Empirical evaluations demonstrate that LPNL achieves significant improvements compared to GNN baselines, showcasing its robust capability in cross-domain knowledge transfer and few-shot learning scenarios.

### 20 Feb 2024 v3

@@ 2.5 Self-Supervised Fine-tuning
As a more relevant graph structure, large-scale graphs lack labelled data. LPNL uses self-supervised learning for large language model fine-tuning. During the end-to-end prompt fine-tuning, it automatically constructs a candidate set containing ground truth, aligned with downstream prediction formats. The ground truth is used as the correct answer for link prediction. To ensure training correctness, the ground truth appears randomly within the candidate neighbor sequence. Notably, during the heterogeneous subgraph sampling process in Sec.2.3, the edges between the ground truth and the source node are masked. Because the self-supervised fine-tuning does not require training labels provided by graph tasks, a fine-tuned LPNL model can make direct predictions on different graphs without the need of extra tuning.
@@ 3.1 Experiment Settings
Models We fine-tune T5-base model Chung et al. (2022) with a 1024 input window constraint as the backbone language model for our LPNL. The numbers of sampling hops $h=2$, top anchor nodes sequence $k=50$, and candidate length limit $L=3$ are used for all following experiments.
Datasets We conducted all experiments on the OAG, known as one of the largest publicly available heterogeneous graphs, comprising 178 million nodes and 2.236 billion edges. It includes five types of nodes (denoted as papers (P), authors (A), venues (V), institutes (I) and fields (F)) and their interrelations. In our specific experiments, we utilized four representative domain-specific subgraphs from OAG: Computer Science (CS), Material Science (Mater), Engineering (Engin) and Chemistry (Chem) Jiang et al. (2021). The graph statistics are listed in Table 1. We partition each dataset into fine-tuning, validation, and test sets based on distinct time periods. Specifically, in the OAG dataset, papers are published between 1900 and 2019. Consequently, we utilize publications preceding 2015 for fine-tuning, data from 2015 to 2016 for validation, and information from 2016 onwards for testing.
Task We consider real-world link prediction tasks to evaluate the performance of our LPNL, specifically, author name disambiguation Ferreira et al. (2012). Author name disambiguation is a fundamental challenge for curating academic publication and author information, as duplicated names are common. The objective is to predict the true author who has a genuine link with a given paper among all authors with the same name.
@@ 3.2 Overall Performance
In this experiment, we compare the T5 model as the backbone version of LPNL to advanced GNN based baseline models across the four domain-specific subgraphs. We fine-tune the model separately across various subgraphs and evaluate the performance of the models in link prediction. The experimental results of the proposed method and baselines are summarized in Table 2. All experiments for the author name disambiguation task over all datasets are evaluated in terms of NDCG, MRR and Hits@1 Li (2022); Liu et al. (2009).
@@
Figure 5: Cross-domain transfer results.
@@ 3.3 Cross-Domain Knowledge Transfer
To explore the generalization capabilities of LPNL, we set up experiments for cross-domain knowledge transfer. Specifically, we fine-tune the T5 model using LPNL on a graph corresponding to one domain and subsequently conducted testing on subgraphs from other domains. The experimental outcomes, visualized in Figure 5 as a heatmap, reveal that in most instances, the model exhibits optimal performance when fine-tuned within its original domain. Surprisingly, the models fine-tuned on other domains also demonstrate remarkably strong performance, often closely matching or even surpassing the best performance achieved by fine-tuning within the original domain (e.g., Mater-Engin). This highlights the robust knowledge transferability of our approach, which means it can make direct predictions on different graphs without the need for additional learning.
Figure 6: LPNL converges fast in few-shot learning compared to GNNs.
3.4 Few-Shot Learning
The extensive pretraining of large language models across various natural language tasks has endowed them with robust reasoning and generalization capabilities. In contrast to traditional GNN models, they require minimal training samples to converge and exhibit superior performance. We further investigates the few-shot learning capabilities of LPNL by comparing it with the top-performing homogeneous GNN and heterogeneous GNN in terms of overall performance. We configure the evaluate results to be printed every 1 batch, with each batch consisting of 50 link prediction tasks. We compare the few-shot results for the first 20 batches. The results in Table 6 demonstrate that our LPNL swiftly converges with minimal sample fine-tuning, displaying comparable performance to the best fine-tuning outcomes. This showcases the portability of large language models in addressing graph-related tasks.
@@ 3.4 Few-Shot Learning
Table 3: Ablation study results of sampling methods.
@@ 3.5 Ablation Study
We conduct an ablation study on CS dataset to evaluate the effectiveness of our approach in employing large language models combined with graph knowledge strategies. We compare the performance among different versions of sampling methods: the standard LPNL, a version without any graph information, and another two sampling versions, each independently utilizing distinct stages. As illustrated in Table 3, the model’s performance significantly diminishes when graph information is excluded. Furthermore, the performance of the versions without stage 1 and stage 2 shows a notable gap compared to LPNL. It indicates that employing our designed two-stage sampling pipeline enables LPNL to capture crucial information within the graph after balanced heterogeneous sampling, resulting in improved predictive outcomes.
 Hop NDCG MRR Hits@1
2-hop 97.86 94.37 89.47
@@
Table 4: Ablation study results of multi-hop sampling.
In our experiments, two critical operations contributing significantly to the outstanding performance of LPNL in link prediction are 2-hop and anchor nodes, which provide essential information to the LLMs. To assess the impact of these two key components on model performance, we conducted another ablation experiments, and the results are presented in Table 4 and 5. It shows that incorporating multi-hop and more anchor nodes information can both enhance the LPNL’s performance. However, further experiments indicate that increasing the number of hops and anchor nodes beyond a certain threshold does not lead to significant performance improvement. On the contrary, it may result in additional costs without notable benefits.
 #Anchor Nodes NDCG MRR Hits@1
@@
Table 5: Ablation study results of top-k anchor nodes
@@ 4 Related Work
Large Language Models with Graph Knowledge The emergence of large language models (LLMs) has propelled natural language processing (NLP) to new heights. Qiu et al. (2020). For example, models like BERT Devlin et al. (2018) and T5 Raffel et al. (2020) demonstrate excellent performance in a wide range of downstream tasks, such as text classification and question answering. Besides, some works  Zhang et al. (2019); Liu et al. (2022, 2020)attempts to inject external graph knowledge into LLMs, thus enabling LLMs to gain the ability to solve problems on graphs. Recently, due to the powerful inferential capabilities of large language models, a burgeoning body of work Fatemi et al. (2023); Ye et al. (2023); Liu et al. (2023) attempt to utilize natural language descriptions of graph features, employing generated prompts to instruct large language models in addressing various problems on graphs.
@@ 5 Discussion
From our experiments, we found that describing graphs using natural language does not follow the principle of "more information is better". Sampling more nodes can introduce additional information, but it may lead to information redundancy, resulting in a decline in the inferential capabilities of large language models. Therefore, the key lies in the setting of the sampling and divide-and-conquer length limits, which should align with the input window size and inferential capabilities of the large language models. While designing prompts, we observe that complex relationships between nodes are challenging to articulate in text, especially in large or dense graphs, potentially leading to redundant contexts. LPNL leverages structural information during the sampling phase and, in the prompt generation, only conveys information about the sampled nodes. This approach aims to minimize context redundancy while maximizing the utilization of graph information.
@@ 6 Conclusion
In this paper, we explore, for the first time, the application of large language models to address the link prediction task on large-scale heterogeneous graphs. We introduce LPNL, a large language models based framework for scalable link prediction on large-scale graphs. We design specific prompt templates for the link prediction task and generate the prompts based on anchor nodes obtained through a two-stage sampling approach. These prompts are then input to the large language models for predictions. To tackle the token overload issue arising from an excessive number of candidate neighbors, we employ a divide-and-conquer strategy. Empirical evaluations demonstrate that LPNL achieves significant improvements compared to GNN baselines, showcasing its robust capability in cross-domain knowledge transfer and few-shot learning scenarios.
Limitations
Some efforts in solving graph-related problems using LLMs involve supervised fine-tuning, resulting in limited ability for knowledge transfer. Although LPNL supports unsupervised learning without the need for labels, it is currently confined to link prediction tasks and has not been applied to a broader spectrum of graph-related tasks. LPNL has not yet explored larger parameter scales for large language models and their zero-shot potentials, which could provide increased input window sizes and enhanced inferential capabilities. Integrating our approach with other graph tasks and larger language models holds the potential to significantly improve predictive capabilities.
Ethics Statement
Ethical considerations are of utmost importance in our research endeavors. In this paper, we conscientiously adhere to ethical principles by exclusively utilizing open-source datasets and employing models that are either open-source or widely recognized in the scientific community. Moreover, our proposed method is designed to ensure that the model does not produce any harmful or misleading information. We are committed to upholding ethical standards throughout the research process, prioritizing transparency, and promoting the responsible use of technology for the betterment of society.